OpenStudy (babyslapmafro):

Two sides of a triangle have lengths of 3cm and 13cm and meet at an angle of 60*. Find the area of the triangle. Thank you...

OpenStudy (campbell_st):

the area of a triangle using trigonometry is  \[A = \frac{1}{2} a \times b \times \sin(C)\] you can use this for your problem